suppose 23% of the individuals in a population have a certain medical condition what is the probability 105 to 135 inclusive will have this in a population of 500
1
Expert's answer
2012-12-04T07:24:25-0500
E(x) = 500 * 0.23 = 115 - the expected result.
P(105 <= x <= 135) = Φ((135 - 115) / 30) - Φ((105 -115) / 30) =
Φ(0.667) - Φ(-0.333) = 0.25 - 0.13 = 12%